OpenStudy (anonymous):

find the volume of pyramid with the hight h and rectangular base with dimension b and 2b

OpenStudy (anonymous):

volume of a pyramid = 1/3 * (height*base area) here base area = b*2b=2b^2 so volume = 1/3 *(2hb^2)

OpenStudy (anonymous):

\[2hb ^{2}/3\]

OpenStudy (anonymous):

dont we need to use integrals?